### v3

# Distributed Algorithms for Potential Problems

Abstract. In this work we present a fast distributed algorithm for local potential problems: these are graph problems where the task is to find a locally optimal solution where no node can unilaterally improve the utility in its local neighborhood by changing its own label. A simple example of such a problem is the task of finding a locally optimal cut, i.e., a cut where for each node at least half of its incident edges are cut edges.
The distributed round complexity of locally optimal cut has been wide open; the problem is known to require $\Omega(\operatorname{log}n)$ rounds in the deterministic LOCAL model and $\Omega(\operatorname{log}\operatorname{log}n)$ rounds in the randomized LOCAL model, but the only known upper bound is the trivial brute-force solution of $O(n)$ rounds. Locally optimal cut in bounded-degree graphs is perhaps the simplest example of a locally checkable labeling problem for which there is still such a large gap between current upper and lower bounds.
We show that in bounded-degree graphs, all local potential problems, including locally optimal cut, can be solved in $\operatorname{log}^{O(1)}n$ rounds, both in the deterministic and randomized LOCAL models. In particular, the deterministic round complexity of the locally optimal cut problem is now settled to $\operatorname{log}^{\Theta(1)}n$.
1 Introduction
In this work, we present a distributed algorithm that finds a locally optimal cut in any bounded-degree graph in a polylogarithmic number of rounds (see Section 1.1); this is the first nontrivial distributed algorithm for this problem. Moreover, our algorithm generalizes to all distributed graph problems that are defined with a local potential function (see Section 1.2).
1.1 Locally optimal cuts
In a graph $G=(V,E)$, a cut is a labeling $\ell:V\to\{-1,+1\}$, an edge $\{u,v\}$ is called a cut edge if $\ell(u)\neq\ell(v)$, and the size of the cut is the total number of cut edges. Maximizing the size of the cut is a classic NP-hard problem.
We say that a cut is locally optimal if for every node at least half of its incident edges are cut edges. Put otherwise, changing the label of $v$ from $\ell(v)$ to $\ell(v)$ does not increase the size of the cut. Such a cut always exists and is trivial to find in the centralized setting: start with any labeling, find any node that violates the constraint (we call these unhappy nodes) and change its label, and repeat until all nodes are happy. Note that changing the label of an unhappy node $v$ will make node $v$ happy but it might make some of its neighbors unhappy. However, this process has to converge after at most $|\absolutevalue{E}$ steps, since the size of the cut increases in each step by at least one.
Distributed graph algorithms.
While it is trivial to find a locally optimal cut with a centralized algorithm, in this work we are interested in distributed algorithms for finding a locally optimal cut. We work in the usual LOCAL model [Lin92, Pel00] of distributed computing: Each node $v\in V$ represents a computer and each edge $\{u,v\}$ represents a communication link between two computers. Initially each node is only aware of its own unique identifier, its degree, and the total number of nodes $n$. Computation proceeds in synchronous rounds; in each round each node can send a message to each neighbor, receive a message from each neighbor, and update its own state. Eventually each node $v$ has to stop and output its own part of the solution, here $\ell(v)$. The running time of the algorithm is $T(n)$ if in any $n$-node graph all nodes stop after at most $T(n)$ rounds.
It is good to note that in this model, time and distance are interchangeable: in $T$ rounds all nodes can gather their radius-$T$ neighborhood, and a $T$-round algorithm is a function that maps radius-$T$ neighborhoods to local outputs. In particular, this means that in $O(n)$ rounds we can solve any graph problem by brute force; the key question is which problems admit a nontrivial $o(n)$-round algorithm.
Distributed algorithms for locally optimal cuts.
It is not hard to turn the trivial centralized algorithm into a valid distributed algorithm for solving the locally optimal cut problem. For example, identify all unhappy nodes, then find a maximal independent set of unhappy nodes, and flip all of them in parallel. Unfortunately, this is not going to be an efficient algorithm. Consider the following graph and a cut $\ell$ indicated with the colors white and black; here the node marked with an arrow is the only unhappy node, as only $1$ of its $3$ incident edges are cut edges:
If we now change the label of this node, we arrive at the following configuration where we again have exactly one unhappy node:
Flipping that we again move unhappiness around:
Clearly this will take $\Omega(n)$ rounds until all unhappiness is resolved (even if we put aside the time needed to find a maximal independent set), so it is no better than the trivial brute-force solution that solves any graph problem in $O(n)$ rounds. Can we do much better than this?
Broader context: locally checkable labelings.
We will study this question in bounded-degree graphs, i.e., we assume that there is some fixed constant $\Delta$ and our goal is to solve locally optimal cut in any graph with maximum degree at most $\Delta$. We note that e.g. in the above apparently-hard instance we had $\Delta=3$, so small values of $\Delta$ do not make the problem trivial.
For any fixed $\Delta$, the locally optimal cut problem is an example of a locally checkable labeling problem (LCL). LCL problems were first defined by [NS95], and there is a large body of work that has developed a complexity theory of LCL problems in the LOCAL model [CV86, Nao91, Lin92, NS95, BFH+16, FG17, GHK18, BHK+18, CP19, CKP19, RG20, BBO+20, BBO+21, DDL+23, Cha20, BBO+21, GRB22, Suo20].
Most natural LCL problems are by now very well-understood, especially for the deterministic LOCAL model. Indeed, for some restricted families of LCLs and restricted families of graphs, there is now a complete classification of the round complexity of all such problems, see e.g. [BBE+20, CSS23, BHK+17, BBC+22, BBC+19]. For example, there is a broad class of symmetry-breaking problems, all of which have round complexity $\Theta(\operatorname{log}^{*}n)$ [CV86, GPS88, Lin92, Nao91], and many other problems [GHK+20] have turned out to be asymptotically as hard as the sinkless orientation problem [BFH+16], which has round complexity $\Theta(\operatorname{log}n)$ in the deterministic LOCAL model [BFH+16, CKP19, GS17].
Locally optimal cut is perhaps the most elementary LCL problem whose distributed complexity is currently wide open. We do not even know if it admits any nontrivial algorithm with round complexity $o(n)$, not even for the smallest nontrivial degree bound $\Delta=3$. (The case of $\Delta=2$ is fully understood, as it is in essence equivalent to the weak-$2$-coloring problem, which can be solved in $O(\operatorname{log}^{*}n)$ rounds, and this is also tight, see e.g. [BHO+19].)
Prior work.
What makes locally optimal cuts particularly intriguing is that many commonly-used lower bound techniques fail. In particular, one cannot use existential graph-theoretic arguments (in the spirit of e.g. the lower bound for coloring trees in [Lin92] or the lower bound for approximate coloring in [CDG+24]), since a locally optimal cut always exists in any graph.
It is, however, possible to construct a nontrivial reduction that connects locally optimal cuts with the sinkless orientation problem [BHL+19], and this gives the following lower bounds: we need at least $\Omega(\operatorname{log}n)$ rounds in the deterministic LOCAL model, and at least $\Omega(\operatorname{log}\operatorname{log}n)$ rounds in the randomized LOCAL model.
This is essentially everything that is known about locally optimal cuts before this work. There are no upper bounds better than the trivial $O(n)$. Furthermore, the lower bounds do not hold in stronger models such as quantum-LOCAL [GKM09, AF14, ACD+25, BBC+25]—there we do not even know if the true complexity of the problem is closer to $O(1)$ or $\Omega(n)$.
We note that locally optimal cuts have been studied e.g. in the context of descriptive combinatorics also under the names of unfriendly colorings and unfriendly partitions [SM12, Con14, CMU20, CT20].
Our contribution.
We present the first nontrivial upper bound for locally optimal cuts. We show that the problem can be solved in $\operatorname{log}^{O(1)}n$ rounds, both in the deterministic and randomized LOCAL models. In particular, the complexity of locally optimal cut in the deterministic LOCAL model is now known to be $\operatorname{log}^{\Theta(1)}n$ rounds.
Open questions.
The main open question is related to models stronger than deterministic LOCAL. In the randomized LOCAL model, there is still an exponential gap between the upper and lower bounds, and in the quantum-LOCAL model we are lacking any nontrivial lower bounds.
1.2 Local potential problems
Local potential problems.
Locally optimal cut is a perfect example of a much more general problem family that we call local potential problems. These are constraint satisfaction problems where we have
  1. 1.
    finitely many possible node labels,
  2. 2.
    a local definition of “happy” vs. “unhappy” nodes,
  3. 3.
    a local potential function that associates a value to each node or edge, and
  4. 4.
    we can turn “unhappy” nodes into “happy” nodes with a local change so that we also locally improve the total potential.
For example, in the locally optimal cut problem, our node labels are $\{-1,+1\}$. A node is “happy” if at least half of its incident edges are cut edges (this is a local definition in the sense that it only depends on the constant-radius neighborhood of the node). We assign local potential $1$ to each cut edge and potential 0 to all other edges (this is a local potential in the sense that it only depends on the constant-radius neighborhood of the edge). Now we can always turn unhappy nodes into happy nodes by flipping their label, and this will improve the total potential around the node.
All local potential problems share the commonality that a feasible solution always exists and it is trivial to find with a centralized algorithm that turns unhappy nodes into happy nodes: as the total potential always strictly increases, and there are only finitely many possible solutions, the centralized algorithm will always converge to an all-happy solution.
Locally optimal problems.
Local potential problems are also closely connected with local optimization problems; these are optimization problems where we have
  1. 1.
    finitely many possible node labels and
  2. 2.
    a local utility function that associates a value to each node or edge,
and the task is to maximize the sum of local utilities. For example, the maximum cut problem is such an optimization problem; the local utility function assigns value $1$ to each cut edge and 0 to all other edges.
Now consider the task of finding a locally optimal solution where we cannot improve the total utility by changing the value of any single node. This is exactly what we have in the locally optimal cut problem. Hence it is an example of finding a locally optimal solution to a local optimization problem, which is a mouthful; we will use the term locally optimal problem for such a task.
It may be helpful to interpret locally optimal problems from a game-theoretic perspective. The goal is to find a stable solution or equilibrium where no node has an incentive to change its label so that it could improve utility in its local neighborhood.
Reductions and generalizations.
Assume we have an efficient algorithm for solving locally optimal problems. Then we can also solve local potential problems equally fast. To see this, note that we can simply ignore the notion of happiness first, and interpret local potentials as utilities. We find a solution where no node can unilaterally improve the values of the local potential functions. Then by definition all nodes will also be happy (as unhappy nodes would have a local move that improves potentials). Hence it is sufficient to design efficient algorithms for locally optimal problems.
At first the definition of locally optimal problems may feel very restrictive, as we only consider solutions where no single node can change its own label to improve the potential. In many problems we would like to make more complicated local changes (e.g. consider the maximum independent set problem, and the task of finding a locally optimal solution where there is no place where removing one node would let us add two nodes). However, in the bounded-degree setting we can easily encode such problems in our formalism. To see this, assume we have a graph problem with $k$ labels. Then the local label of a node $v$ can be a vector of $k^{\Delta+1}$ labels $\ell_{0},\dotsc,\ell_{\Delta}$, where $\ell_{0}$ is the “base” label of node $v$, and $\ell_{i}$ indicates that node $v$ wants to add $\ell_{i}$ modulo $k$ to the label of its $i$th neighbor. This way changing the physical label of one node allows us to change the logical labels of all nodes in its radius-$1$ neighborhood. This scheme of course generalizes also to any constant radius. Hence without loss of generality it suffices to consider solutions where no single node can unilaterally improve the total utility.
Our contribution.
In this work we design an algorithm that solves any locally optimal problem in $\operatorname{log}^{O(1)}n$ rounds in bounded-degree graphs in the deterministic LOCAL model. As a corollary, we can also solve any local potential problem in the same time. Our result for locally optimal cuts now follows as a special case of this.
Our work also makes it possible to directly translate many centralized optimization algorithms that are based on local search into nontrivial distributed algorithms. Our work also gives a new, simple way to argue that a given LCL problem $\Pi$ can be solved in polylogarithmic time: simply try to come up with a local potential function or a local utility function that allows us to interpret $\Pi$ as a local potential problem or as a locally optimal problem.
Conversely, super-polylogarithmic lower bounds on the distributed round complexity now directly imply that a problem cannot be represented with the help of a local potential. To give a simple example, stable matchings require a linear number of rounds, even in bounded-degree bipartite graphs [FKP+10], so this problem cannot be represented with local potential functions. Results of this form may be of interest also in a much broader context outside distributed computing.
An application: defective coloring.
Key examples of local potential problems include defective coloring problems, see e.g. [BE13, Chapter 6], [CPS17, FG17]. Assume that we have $c$ colors and we want to have defect at most $d$, that is, each node can have at most $d$ neighbors with the same color. If $c(d+1)>\Delta$, this is a local potential problem. Here the potential function is the number of properly colored edges. If a node is unhappy (it has more than $d$ neighbors with the same color), it can find a color that makes it happy. Our algorithm can be used to solve this problem in for any constants $c$, $d$ and $\Delta$ with $c(d+1)>\Delta$ in polylogarithmic number of rounds.
Future work and open questions.
We believe the notions of local potential problems and locally optimal problems will inspire plenty of follow-up work; this can also serve as a bridge between distributed constraint satisfaction problems (e.g. LCL) and distributed optimization and approximation problems.
A key open question is understanding the distributed complexity landscape of the family of local potential problems. Our work together shows that there are local potential problems with deterministic round complexity $\operatorname{log}^{\Theta(1)}n$, and by prior work there are also local potential problems with round complexity $\Theta(\operatorname{log}^{*}n)$ and $O(1)$. However, one question that is currently wide open is what happens when we step outside the deterministic LOCAL model. For example, does randomness ever help with any of these problems? The only lower bound we have is the exponentially lower $\Omega(\operatorname{log}\operatorname{log}n)$ for locally optimal cuts, is this the correct complexity? If yes do all local potential problems admit $\operatorname{poly}(\operatorname{log}\operatorname{log}n)$ randomized algorithms? Do any of these problems admit a distributed quantum advantage?
Also questions related to decidability are wide open. Fore example, given the specification of a local potential problem, is it decidable whether it can be solved in $O(\operatorname{log}^{*}n)$ rounds or whether it requires $\Omega(\operatorname{log}n)$ rounds?
Finally, our work only considered the case of bounded-degree graphs. Can we solve local potential problems in polylogarithmic time also in general graphs?
1.3 New techniques and key ideas
Informal intuition.
Our algorithm is based on the idea that we start with an arbitrary labeling and repeatedly improve it. However, if we only fix nodes that are unhappy, we will end up in a long dependency chain, similar to what we discussed in Section 1.1.
Intuitively, the key challenge in our example was that the solution was very fragile—all nodes were barely happy, and a minor perturbation in their local neighborhood may result them to flip, which leads to long cascades that we must avoid.
What our algorithm does differently is that we push it further, beyond the minimal requirement of happiness. Our algorithm tries to find local neighborhoods where the value of the potential can be improved by changing multiple nodes. For example, the algorithm might notice that the shaded area here admits a modification that improves the total potential a lot:
Such a change may at first look counterproductive, as we did not make any nodes happy, yet we increased the number of unhappy nodes. However, if we now start fixing unhappy nodes near the left end of the graph, such a fixing process will not propagate deep inside the shaded region. Intuitively, the nodes in the shaded region are so happy that minor perturbations will not result in catastrophic cascading chains of correction.
Key properties (Section 4).
Let us now try to identify a key combinatorial property of locally optimal cuts (and more generally, local potential problems). Consider first a sequential algorithm that eliminates all unhappy nodes one by one, let $X$ be the set of all nodes that were flipped in this process at least once. Note that in general, $X$ might span the entire input graph.
We prove that for every node $u\in X$, we can find in the $O(\operatorname{log}n)$-radius neighborhood of $u$ a set of nodes $A$ that we can flip so that we get a significant improvement in the total potential. For example, in the above illustration, the changes that we did in the shaded area improved the total potential significantly.
Conversely, this means that if there is no such improving set $A$ in the $O(\operatorname{log}n)$-radius neighborhood of node $u$, we know that we are safe: the sequential flipping procedure will never reach node $u$.
The goal of the algorithm is to reach a situation in which all nodes can declare that they are safe. Then they are not only currently happy, but also cannot be influenced by any future changes.
Algorithm (Section 3).
Given the above property, the key challenge is to orchestrate the process of making all nodes safe in a parallel and distributed manner.
Our algorithm uses the MPX clustering algorithm [MPX13, EN22] to divide the nodes into polylogarithmic-radius clusters with small boundaries. Each cluster can make nodes far from boundaries safe; we can simply compute for each cluster the result of repeatedly finding for each internal node an improving set $A$, and flipping $A$ until all internal nodes are safe.
The main technical challenge is taking care of nodes close to boundaries that may not be safe and not even happy. To this end, we repeat the randomized clustering process $O(\operatorname{log}n)$ times, to ensure that each node will be w.h.p. at least once far from a boundary and hence made safe.
Now we need to ensure that we do not undo what we gained in the previous phase. In particular, consider a node $u$ that was deep inside a cluster in phase $1$, and was made safe, but it is close to a cluster boundary in phase $2$. We need to ensure that improvements in phase $2$ do not tamper with the property that node $u$ was already safe.
Here we make use of the idea that we can have different “degrees” of safety: In phase $1$, we are more aggressive in making nodes safe than in phase $2$; put otherwise, in phase $2$ we require a larger relative gain when we are finding improving sets.
This is sufficient to ensure that node $u$ is protected from a chain of improvements in phase $2$. To see this, we consider a sequence of improvements that we did in phase $2$ in the vicinity of node $u$. The union of these strongly improving sets we found in phase $2$ would give a weakly improving set that we would have found already in phase $1$ around node $u$, but as $u$ was made very safe, such a set cannot exist. Hence phase-$2$ improvements will not tamper with phase-$1$ gains.
A careful choice of parameters allows us to push this reasoning throughout all $O(\operatorname{log}n)$ phases, showing that each node is made safe and hence in particular happy during the process, and this property is never lost.
Our algorithm is randomized, but it can be derandomized (with polylogarithmic cost) by using the standard techniques [GHK18, RG20].
Notes on the definitions (Section 2).
So far in the introduction, we have used the locally optimal cut problem as a running example, and there it is natural to relate it with the max-cut problem and use the number of cut edges as the potential or utility function; in particular, unhappy nodes can be flipped to increase the potential. In the technical parts, it is more convenient to use a convention where unhappy nodes can be flipped to decrease the potential. Hence the formal definition of potential problems in Section 2 follows this convention.
2 Preliminaries
In this section, we provide some basic graph definitions, and then we define the family of problems that we consider.
2.1 Notation and definitions
In this paper, we consider simple undirected graphs unless otherwise stated. For a graph $G=(V,E)$, we denote its set of nodes by $V$ and its set of edges by $E$. If $G$ is not clear from the context we use $V(G)$ to denote the set of nodes of $G$ and $E(G)$ to denote the set of edges of $G$. For any two nodes $u,v\in V(G)$ the distance between $u$ and $v$ is the number of hops in a shortest path between $u$ and $v$, and is denoted by $\operatorname{dist}(u,v)$. For any node $v\in V(G)$, the radius-$r$ neighborhood $\mathcal{N}_{r}[v]$ of $v$ is the set of all nodes at distance at most $r$ from $v$. Let $A\subseteq V(G)$ be a subset of nodes, and $B\subseteq E(G)$ a subset of edges. By $\mathcal{N}_{r}[A]$ we denote the set of all nodes at distance at most $r$ from any node in $A$, and by $\mathcal{N}_{r}[B]$ we denote the set of all nodes that are endpoints of edges in $B$ or are at distance at most $r$ from an endpoint of an edge in $B$. For any two graphs $G,H$, the intersection between $G$ and $H$ is the graph $G\cap H=(V(G)\cap V(H),E(G)\cap E(H))$, and the union between $G$ and $H$ is the graph $G\cup H=(V(G)\cup V(H),E(G)\cup E(H))$. The difference between two graphs $G$ and $H$ is the graph $G\setminus H=(V(G),E(G)\setminus E(H))$.
For any graph $G=(V,E)$ and any node $v\in V$, if $\{v,u\}\in E$ is an edge incident to $v$, we say that the pair $(v,\{v,u\})$ is a half-edge incident to $v$. We denote the set of all half-edges in the graph $G$ by $\mathcal{H}(G)$. By $[x]$ we denote the set $\{1,\ldots,x\}$.
Definition 2.1 (Labeled graph).
Let $\mathcal{V}$ and $\mathcal{E}$ be two sets of labels. We say that a graph $G=(V,E)$ is $(\mathcal{V},\mathcal{E})$-labeled if there exist two labeling functions $\ell_{\operatorname{nodes}}\colon V\to\mathcal{V}$ and $\ell_{\operatorname{half-edges}}\colon\mathcal{H}(G)\to\mathcal{E}$ that assign labels to the nodes and half-edges of $G$, respectively. The pair $\ell=(\ell_{\operatorname{nodes}},\ell_{\operatorname{half-edges}})$ is said to be the labeling of $G$.
LOCAL model.
In this paper, we consider the LOCAL model of distributed computation [Lin92]. We are given a distributed network of $n$ nodes/processors, each of which is capable of arbitrary (but finite) computation. Nodes are connected by edges which represent communication links, and are able to send messages of arbitrarily large (but finite) size to each other: note that nodes can discriminate between different communication links through ports, which are numbered from $1$ to $\Delta$, where $\Delta$ is the maximum degree of the network, that are adversarially assigned. Each node is also given a unique identifier in the set $n^{c}]$ for some constant $c\geq 1$, which is, again, adversarially assigned. The communication network is represented by a simple undirected graph $G=(V,E)$ where $V$ is the set of nodes/processors and $E$ is the set of edges/communication links. All nodes run the same algorithm and computation proceeds in synchronous rounds. In each round, each node sends messages to its neighbors, receives messages from all of its neighbors, and performs local computation. The input of a problem is given in the form of a $(\mathcal{V},\mathcal{E})$-labeled graph $G=(V,E)$ where $\mathcal{V}$ and $\mathcal{E}$ are sets of labels. The output of a node is an assignment of labels to the node itself and/or to its incident half-edges: note that this information can be stored in the node through the use of port numbers. When a node determines its own output, we say that the node terminates its computation. The running time of a LOCAL algorithm is the number of communication rounds it takes for all nodes to produce their output. We also refer to the running time of an algorithm as the locality of the algorithm. The complexity of a problem is the minimum running time (across all possible algorithms) needed to solve the problem. We can also assume that nodes are equipped with some source of randomness (independently of other nodes), such as an arbitrarily long random bit string, which they can use during the computation. In this case, we talk about the randomized LOCAL model as opposed to the deterministic LOCAL model. In case of randomized computation, we require that the problem is solved with probability at least 11/n1-1/n.
2.2 Problems of interest
We first define the notion of centered graph, that is, a graph with a distinguished node that we call the center.
Definition 2.2 (Centered graph).
Let $H=(V,E)$ be any graph, and let $v\in V$ be any node of $H$. The pair $(H,v)$ is said to be a centered graph and $v$ is said to be the center of $(H,v)$. We say that the pair $(H,v)$ has radius $r$ if all nodes in $H$ are at distance at most $r$ from $v$ and all edges in $H$ have at least one endpoint at distance at most r1r-1 from $v$.
Definition 2.3 (Labeled centered graph).
Let $\mathcal{V},\mathcal{E}$ be two finite sets of labels, and $r,\Delta\in\mathbb{N}$. We denote by $\mathcal{L}(\mathcal{V},\mathcal{E},r,\Delta)$ the family of all centered graphs $(H,v)$ such that $H$ is a $(\mathcal{V},\mathcal{E})$-labeled graph of degree at most $\Delta$ and has radius at most $r$.
We will focus on problems that ask to produce labelings over graphs that satisfy some constraints in all radius-$r$ neighborhoods of every node. Consider a centered graph $(H,v)$.
Definition 2.4 (Set of constraints).
Let $r,\Delta\in\mathbb{N}$ be constants, $I$ a finite set of integers, and $(\mathcal{V},\mathcal{E})$ a tuple of finite label sets. Let $\mathcal{C}$ be a finite set of centered graphs $\{(H_{i},v_{i})\}_{i\in I}$ where each $H_{i}$ is a $(\mathcal{V},\mathcal{E})$-labeled graph of degree at most $\Delta$ and $(H_{i},v_{i})$ has radius at most $r$ for all $i\in I$. Then, $\mathcal{C}$ is said to be an $(r,\Delta)$-set of constraints over $(\mathcal{V},\mathcal{E})$.
In the following, we define what it means to satisfy a set of constraints. For a $(\mathcal{V},\mathcal{E})$-labeled graph $G=(V,E)$ and a node $v\in V$, we denote by $G_{r}(v)$ the $(\mathcal{V},\mathcal{E})$-labeled subgraph of $G$ that contains all nodes at distance at most $r$ from $v$ and all edges that have at least one endpoint at distance at most r1r-1 from $v$.
Definition 2.5 (Labeled graph satysfying a set of constraints).
Let $G=(V,E)$ be a $(\mathcal{V},\mathcal{E})$-labeled graph for some finite sets $\mathcal{V}$ and $\mathcal{E}$ of labels. Let $\mathcal{C}$ be an $(r,\Delta)$-set of constraints over $(\mathcal{V},\mathcal{E})$ for some finite $r,\Delta\in\mathbb{N}$. We say that the graph $G$ satisfies $\mathcal{C}$ if, for every node $v\in V$, it holds that $(G_{r}(v),v)\in\mathcal{C}$.
Before introducing the family of problems considered in this paper, we first define the family of problems called Locally Checkable Labelings.
Definition 2.6 (Locally checkable labeling (LCL) problems).
Let $r,\Delta\in\mathbb{N}$ be constants. A locally checkable labeling (LCL) problem $\Pi$ is a tuple $(\mathcal{V}_{\operatorname{in}},\mathcal{E}_{\operatorname{in}},\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{out}},\mathcal{C})$ where $\mathcal{V}_{\operatorname{in}}$, $\mathcal{E}_{\operatorname{in}}$, $\mathcal{V}_{\operatorname{out}}$, and $\mathcal{E}_{\operatorname{out}}$ are finite label sets and $\mathcal{C}$ is an $(r,\Delta)$-set of constraints over $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$.
Solving an LCL problem $\Pi$ means the following. We are given as input a $(\mathcal{V}_{\operatorname{in}},\mathcal{E}_{\operatorname{in}})$-labeled graph $G=(V,E)$. For any node $v\in V$ and half-edge $(v,e)\in\mathcal{H}(G)$, let us denote the input label of $v$ by $\ell_{\operatorname{in}}(v)$, and the input label of $(v,e)$ by $\ell_{\operatorname{in}}((v,e))$. We want to produce an output labeling on $G$ so that we obtain a $(\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{out}})$-labeled graph. Let us denote by $\ell_{\operatorname{out}}(v)$ and by $\ell_{\operatorname{out}}((v,e))$ the output labels of $v$ and $(v,e)$, respectively. Consider the $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph $G$ where the labeling is defined as follows: the label of each node $v$ is $\ell_{\operatorname{nodes}}(v)=(\ell_{\operatorname{in}}(v),\ell_{\operatorname{out}}(v))$, and the label of each half-edge $(v,e)$ is $\ell_{\operatorname{half-edges}}((v,e))=(\ell_{\operatorname{in}}((v,e)),\ell_{\operatorname{out}}((v,e)))$. Now, the $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph $G$ must satisfy the $(r,\Delta)$-set of constraints $\mathcal{C}$ over $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$ according to Definition 2.5.
With these notions in mind, we can define the family of Locally Optimal Problems (LOPs) that we consider in this paper.
Definition 2.7 (Locally Optimal Problems).
A Locally Optimal Problem (LOP) is a pair $(\Pi,\Psi)$ where:
  • $\Pi=(\mathcal{V}_{\operatorname{in}},\mathcal{E}_{\operatorname{in}},\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{out}},\mathcal{C})$ is an LCL problem as defined in Definition 2.6, where $\mathcal{C}$ is a $(2r,\Delta)$-set of constraints over $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$.
  • $\Psi\colon\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},r,\Delta)\to\mathbb{R}_{\geq 0}$ is a function that assigns a non-negative real value to every $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled centered graph $(H,v)$.
Moreover, $\mathcal{C}$ must satisfy the following property, which in essence says that a correct solution does not allow local improvements. For any $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph $G=(V,E)$, let us denote by $\ell=(\ell_{\operatorname{nodes}},\ell_{\operatorname{half-edges}})$ the labeling of $G$. We say that the potential associated to $(G,\ell)$ is $\operatorname{Pot}(G,\ell)=\sum_{v\in V}\Psi_{\ell}(G_{r}(v),v)$, where $\Psi_{\ell}(G_{r}(v),v)$ is the evaluation of $\Psi$ on $(G_{r}(v),v)$ that is labeled by $\ell$. For any node $v\in V$, let the cost of $v$ be $c_{v}:=\Psi_{\ell}(G_{r}(v),v)$. Then, the centered graph $(G_{2r}(v),v)$ labeled with $\ell$ belongs to $\mathcal{C}$ if and only if we cannot change the labeling of $v$ to improve the potential of a node in $(G_{2r}(v),v)$. Formally, there is no labeling $\ell^{\prime}$ with the following properties, where, for any node $u\in V$, $c_{u}^{\prime}:=\Psi_{\ell^{\prime}}(G_{r}(u),u)$.
  • $\ell^{\prime}$ differs from $\ell$ only on the output labels of $v$ and of its incident half-edges.
  • According to the labeling $\ell^{\prime}$, $G_{2r}(v)$ is still a $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph.
  • It holds that $c_{v}^{\prime}<c_{v}$, and that $\operatorname{Pot}(G_{r}(v),\ell^{\prime})<\operatorname{Pot}(G_{r}(v),\ell)$.
In other words, a centered graph $(G_{2r}(v),v)$ is contained in $\mathcal{C}$ if and only if there is no alternative labeling for $v$ where:
  • $\Psi$ gives a strictly better value on node $v$, and
  • the overall potential is strictly better (observe that the changes on $v$ can only affect nodes within distance $r$ from $v$, and their neighborhood is contained in the radius-$2r$ neighborhood of $v$).
The reader might wonder why we do not consider a slightly broader definition of LOPs, where we allow the relabeling to take place on all nodes and half-edges of $G_{r}(v)$. In Definition 2.8 we provide such a more general definition, and in Lemma 2.9 we will show that, given an algorithm for solving LOPs, we can use such an algorithm to solve problems in this broader family.
Definition 2.8 (Generalized Locally Optimal Problems).
A Generalized Locally Optimal Problem (GLOP) is a pair $(\Pi,\Psi)$ where:
  • $\Pi=(\mathcal{V}_{\operatorname{in}},\mathcal{E}_{\operatorname{in}},\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{out}},\mathcal{C})$ is an LCL problem as defined in Definition 2.6, where $\mathcal{C}$ is a $(3r,\Delta)$-set of constraints over $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$.
  • $\Psi\colon\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},r,\Delta)\to\mathbb{R}_{\geq 0}$ is a function that assigns a non-negative real value to every $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled centered graph $(H,v)$.
Moreover, $\mathcal{C}$ must satisfy the following property. For any $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph $G=(V,E)$, let us denote by $\ell=(\ell_{\operatorname{nodes}},\ell_{\operatorname{half-edges}})$ the labeling of $G$. We say that the potential associated to $(G,\ell)$ is $\operatorname{Pot}(G,\ell)=\sum_{v\in V}\Psi_{\ell}(G_{r}(v),v)$, where $\Psi_{\ell}(G_{r}(v),v)$ is the evaluation of $\Psi$ on $(G_{r}(v),v)$ that is labeled by $\ell$. The additional requirement that $\mathcal{C}$ needs to satisfy is that, if the centered graph $(G_{3r}(v),v)$ labeled with $\ell$ does not belong to $\mathcal{C}$, then there must exist a labeling $\ell^{\prime}$ with the following properties.
  • Labeling $\ell^{\prime}$ differs from $\ell$ only on the output labels of $G_{r}(v)$.
  • According to the labeling $\ell^{\prime}$, $G_{3r}(v)$ is still a $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph.
  • The centered graph $(G_{r}(v),v)$ labeled with $\ell^{\prime}$ belongs to $\mathcal{C}$.
  • It holds that $\operatorname{Pot}(G_{2r}(v),\ell^{\prime})<\operatorname{Pot}(G_{2r}(v),\ell)$.
In other words, if a centered graph $(G_{3r}(v),v)$ labeled with $\ell$ is not contained in $\mathcal{C}$, then the outputs on $(G_{r}(v),v)$ can be changed such that we obtain a centered graph in $\mathcal{C}$, and the obtained potential is strictly better. (Observe that the changes on $G_{r}(v)$ can only affect nodes within distance $2r$ from $v$, and their neighborhood is contained in the radius-$3r$ neighborhood of $v$.)
We now prove that we can use an algorithm designed to solve LOPs to solve GLOPs.
Lemma 2.9.
Consider a GLOP $(\Pi,\Psi)$ as defined in Definition 2.8. Then, there exists an LOP $(\Pi^{\prime},\Psi^{\prime})$ as defined in Definition 2.7 such that any solution to $(\Pi^{\prime},\Psi^{\prime})$ can be converted into a solution of $(\Pi,\Psi)$ by an $O(1)$-round LOCAL algorithm.
Proof.
Let $(\Pi,\Psi)$ be a GLOP as defined in Definition 2.8, where the constraint of $\Pi$ contains centered neighborhoods of radius $3r$. In the following, we define $(\Pi^{\prime},\Psi^{\prime})$. We start by defining the labels of $\Pi^{\prime}$.
  • The sets of input labels is the same, i.e., $\mathcal{V}_{\operatorname{in}}^{\prime}=\mathcal{V}_{\operatorname{in}}$, $\mathcal{E}_{\operatorname{in}}^{\prime}=\mathcal{E}_{\operatorname{in}}$.
  • Let $d_{r,\Delta}$ be an upper bound on the number of nodes and half-edges contained in a centered graph of maximum degree $\Delta$ and radius $r$. The sets of output labels are as follows.
    • $\mathcal{V}_{\operatorname{out}}^{\prime}=\{\bot\}$.
    • $\mathcal{E}_{\operatorname{out}}^{\prime}=[\Delta]\times[|\mathcal{V}_{\operatorname{out}}|]^{d_{r,\Delta}}\times[|\mathcal{E}_{\operatorname{out}}|]^{d_{r,\Delta}}$.
  • The radius of the centered graphs in $\mathcal{C}^{\prime}$ is $3r^{\prime}=6r$ and the maximum degree is $\Delta^{\prime}=\Delta$.
For any $(\mathcal{V}_{\operatorname{in}}^{\prime}\times\mathcal{V}_{\operatorname{out}}^{\prime},\mathcal{E}_{\operatorname{in}}^{\prime}\times\mathcal{E}_{\operatorname{out}}^{\prime})$-labeled graph $G=(V,E)$, let us denote by $\ell=(\ell_{\operatorname{nodes}},\ell_{\operatorname{half-edges}})$ the labeling of $G$. For each node $v$, we now define a function $f$ that takes as input a centered graph $(G_{r}(v),v)$ labeled with $\ell$, and produces a labeling. By abusing the notation, with $f$ we also denote the obtained labeling. The function $f$ will be a partial function, that is, in some cases, the function $f$ will not be able to label $(G_{r}(v),v)$.
Let $(p_{1},X_{1},Y_{1}),\ldots,(p_{\deg(v)},X_{\deg(v)},Y_{\deg(v)})$ be the labels assigned by $\ell$ to the half-edges incident to $v$. The values $p_{1},\ldots,p_{\deg(v)}$ are called ports of $v$. If $\exists i,j~:~X_{i}\neq X_{j}\text{ or }Y_{i}\neq Y_{j}$, then $f((G_{r}(v),v))$ is undefined. Let $X=X_{1}$ and $Y=Y_{1}$. Consider a BFS visit on $G_{r}(v)$ where, the order in which the neighbors of some node $u$ are visited, is according to the ports of $u$. If $u$ has ports with values not in $\deg(u)]$, or two incident ports with the same value, then $f$ is undefined. The labeling $f$ on $(G_{r}(v),v)$ is then obtained by assigning, to the $j$-th visited half-edge, the $j$-th element of $X$, and to the $j$-th visited node, the $j$-th element of $Y$.
We now define a function $g$ that takes as input a centered graph $(G_{2r}(v),v)$ labeled with $\ell$ and returns a labeling for $(G_{r}(v),v)$ that uses the labels of $\Pi$. By abusing the notation, with $g$ we also denote the obtained labeling. Again, the function $g$ will be a partial function. For each node $u$ in $(G_{r}(v),v)$, compute $f((G_{r}(u),u))$. If $f$ is undefined in any of the cases in which we applied it, then $g$ is undefined on $(G_{2r}(v),v)$. For each node (resp. half-edge) $z$ in $(G_{r}(v),v)$, the labeling $g$ for $\Pi$ is defined as follows. Let $a$ be the array containing the output that each node in $(G_{r}(v),v)$ produced for $z$, according to $f$. Then, $g$ labels $z$ with the $(\sum_{x\in a}x\pmod{|\mathcal{V}_{\operatorname{out}}|})-$th label in $\mathcal{V}_{\operatorname{out}}$ (resp. $(\sum_{x\in a}x\pmod{|\mathcal{E}_{\operatorname{out}}|})-$th label in $\mathcal{E}_{\operatorname{out}}$), according to some arbitrary but fixed order.
The potential function $\Psi^{\prime}$ according to $\ell$ is defined as follows.
$\Psi^{\prime}_{\ell}(G_{r^{\prime}}(v),v):=\Psi_{g}(G_{r}(v),v)\text{ if $g$ on $(G_{r^{\prime}=2r}(v),v)$ is defined, and}$
$\Psi^{\prime}_{\ell}(G_{r^{\prime}}(v),v):=M\text{ otherwise,}$
where $M$ is an arbitrary value strictly larger than any value returned by $\Psi$.
The constraint $\mathcal{C}^{\prime}$ contains exactly those $\ell$-labeled centered graphs $(G_{3r^{\prime}}(v),v)$ such that there is no labeling $\ell^{\prime}$, that differs only on the output labels of $v$ and its incident half-edges, that satisfies $\Psi^{\prime}_{\ell^{\prime}}(G_{r^{\prime}}(v),v)<\Psi^{\prime}_{\ell}(G_{r^{\prime}}(v),v)$ and $\operatorname{Pot}(G_{r^{\prime}}(v),\ell^{\prime})<\operatorname{Pot}(G_{r^{\prime}}(v),\ell)$. This condition ensures that the LOP $(\Pi^{\prime},\Psi^{\prime})$ that we defined satisfies the requirements of Definition 2.7.
Now, we prove that any solution for $(\Pi^{\prime},\Psi^{\prime})$ can be converted into a solution for $(\Pi,\Psi)$ in $O(1)$ rounds. We first observe that, in any valid solution for $(\Pi^{\prime},\Psi^{\prime})$, the potential on each node $v$ must be strictly less than $M$. In fact, if the potential computed on node $v$ is $M$, it means that $g$ is undefined on the $r^{\prime}=2r$-radius neighborhood of $v$. This implies that $f$ is undefined on some node $u$ within distance $r$ from $v$, and the potential on $u$ is hence also $M$. Observe that, in this case, $u$ can improve its potential by producing an output that encodes a valid port assignment of its $r$-hop neighborhood. ∎
3 The algorithm
3.1 Preliminary definitions and results
A key notion in our algorithm is that of an improving set. As already described in the introduction, the main insight of our approach is that if a node $u$ is part of some sequential flipping process that obtains a globally correct solution, we can detect this in $u$’s $O(\operatorname{log}n)$ neighborhood. What we will find in $u$’s neighborhood is an improving set with a good improving ratio. We will later prove this formally, so to state our algorithm, we give the necessary definitions here.
Definition 3.1 (Improving set).
Let $(\Pi,\Psi)$ be an LOP as defined in Definition 2.7. Consider any $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph $G=(V,E)$ not necessarily satisfying the constraints of $\Pi$. Let $\ell_{1}$ denote the labeling of $G$. For a subset $A\subseteq V$ of nodes, we denote by relabeling of $A$ a new output label assignment to the nodes of $A$ and to their incident half-edges. Consider any subset $A$ that can be relabeled, obtaining a new labeling $\ell_{2}$ of $G$, such that $\operatorname{Pot}(G,\ell_{2})<\operatorname{Pot}(G,\ell_{1})$. We say that the pair $(A,\ell_{2})$ is an improving set of $G$ with respect to $\ell_{1}$. The improvement of $(A,\ell_{2})$ with respect to $\ell_{1}$ is defined as the difference $\operatorname{Imp}(A,\ell_{1},\ell_{2}):=\operatorname{Pot}(G,\ell_{1})-\operatorname{Pot}(G,\ell_{2})$. The improving ratio (IR) of $(A,\ell_{2})$ with respect to $\ell_{1}$ is defined as the ratio $|\operatorname{IR}(A,\ell_{1},\ell_{2}):=\operatorname{Imp}(A,\ell_{1},\ell_{2})/\absolutevalue{A}$ if $|A|>0$. For an empty set $A$ of nodes we define $\operatorname{IR}(A,\ell_{1},\ell_{2}):=0$, for any $\ell_{1},\ell_{2}$. By diameter of an improving set $(A,\ell_{A})$, we denote the weak diameter of the subgraph $G[A]$ induced by $A$.
Definition 3.2 (Minimal improving set).
An improving set $(A,\ell_{2})$ is said to be minimal if there is no improving set $(A^{\prime},\ell_{2}^{\prime})$ such that $A^{\prime}\subseteq A$ and $\operatorname{IR}(A^{\prime},\ell_{1},\ell_{2}^{\prime})>\operatorname{IR}(A,\ell_{1},\ell_{2})$.
Notice that every improving set $(A,\ell_{2})$ contains a minimal improving set, as one can iteratively pick a subset of nodes (possibly, the same set) and a new labeling that increases the improving ratio until no such subset exists. Note that this process must terminate with at least one node, as the empty graph has improving ratio 0.
From a top-down view, our algorithm does nothing else than relabeling a lot of improving sets. So the changes that our algorithm does to the labeling can be fully described as a sequence of improving sets.
Definition 3.3 (Sequence of $\beta$-improving sets).
Let $(\Pi,\Psi)$ be an LOP. Consider any $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph $G=(V,E)$ not necessarily satisfying the constraints of $\Pi$, and let $\ell_{0}$ denote the labeling of $G$. Let $\beta>0$. A sequence of $\beta$-improving sets in $G$ w.r.t. $\ell_{0}$ is a sequence of pairs $(A_{1},\ell_{1}),(A_{2},\ell_{2}),\ldots,(A_{k},\ell_{k})$ such that:
  • For every $1\leq i\leq k$, the pair $(A_{i},\ell_{i})$ is a minimal improving set of $G$ w.r.t. the labeling $\ell_{i-1}$.
  • For every $1\leq i\leq k$, it holds that $\operatorname{IR}(A_{i},\ell_{i-1},\ell_{i})\geq\beta$.
To find improving sets and figure out which improving sets to relabel, our algorithm will first compute a clustering of the graph and then find a good sequence of improving sets in every cluster. Note that finding such a sequence inside a cluster is easy if the cluster has small diameter, as we can just brute force this computation.
MPX clustering.
In this paragraph, we define and describe the MPX clustering algorithm, which is a key component of our algorithm, following the paper that first introduced it [MPX13].
Let us define the notion of a $(\rho,d)$-decomposition of a graph $G=(V,E)$.
Definition 3.4 ($(\rho,d)$-decomposition of a graph).
A $(\rho,d)$-decomposition of a graph $G=(V,E)$ is a partition of the nodes $V$ into sets $C_{1},\ldots,C_{k}$, which we call clusters, such that:
  • For every $i$, the strong diameter of each $C_{i}$ is at most $d$.
  • The number of edges whose endpoints belong to different clusters is at most $\rho|E|$.
Before stating the properties of the MPX clustering algorithm, we report a useful lemma.
Lemma 3.5 (Lemma 4.4 of [MPX13]).
Let $d_{1}\leq\ldots\leq d_{n}$ be arbitrary values, and let $\lambda_{1},\ldots,\lambda_{n}$ be independent random variables such that $\lambda_{i}\sim\text{Exp}(\beta)$ for each $i$, for some $\beta>0$. Then, the probability that the smallest and the second-smallest values of $d_{i}-\lambda_{i}$ are within distance $c$ of each other is at most $O(\beta c)$.
The MPX clustering algorithm achieves the following result.
Lemma 3.6 ([MPX13]).
There exists a randomized LOCAL algorithm that computes a $(\rho,d)$-decomposition of a graph $G=(V,E)$ of $n$ nodes in $O(\operatorname{log}n/\rho)$ rounds w.h.p., for some $d=O(\operatorname{log}n/\rho)$. Let $k=1/(c\rho)$, where $c$ is a large enough constant. Furthermore, it holds that for each $v\in V$, the probability that the radius-$k$ neighborhood of $v$ is fully contained in some cluster is at least $1/2$.
Proof.
We briefly show how the MPX algorithm works. Every node $u$ of the input graph $G=(V,E)$ is assigned, independently of other nodes, an exponential random variable $\lambda_{u}\sim\text{Exp}(\rho/2)$. We say that $\lambda_{u}$ is the shift of $u$. The shifting distance from a node $u$ to a node $v$ is defined as $\operatorname{dist}_{\text{shift}}(u,v)=\operatorname{dist}(u,v)-\lambda_{u}$: note that, with probability 1, $\operatorname{dist}_{\text{shift}}(u,v)\neq\operatorname{dist}_{\text{shift}}(v,u)$. Each node $u$ that, for any other node $v\neq u$, satisfies $\operatorname{dist}_{\text{shift}}(u,u)<\operatorname{dist}_{\text{shift}}(v,u)$, will be the center of the cluster $C_{u}$. We show that at least one such node exists.
Consider any node $v\in V$, and take the node $u$ that realizes $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$. By contradiction, suppose there exists a node $w$ such that $\operatorname{dist}_{\text{shift}}(u,u)>\operatorname{dist}_{\text{shift}}(w,u)$. Then, we have that $\lambda_{u}>\operatorname{dist}(w,u)-\lambda_{w}$. This implies that $\operatorname{dist}_{\text{shift}}(u,v)=\operatorname{dist}(u,v)-\lambda_{u}>\operatorname{dist}(u,v)+\operatorname{dist}(u,w)-\lambda_{w}\geq\operatorname{dist}(w,v)-\lambda_{w}=\operatorname{dist}_{\text{shift}}(w,v)$ by the triangular inequality. Hence, $u$ does not realize the minimum of $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$, reaching a contradiction.
Now, we say that each node $v$ joins the cluster of the node $u$ that minimizes the shifting distance from $u$ to $v$, that is, $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$. Let $d=O(\operatorname{log}n/\rho)$. In [MPX13], it is shown that this procedure produces a $(\rho,d)$-decomposition of $G$ with high probability. While the original procedure is defined for the PRAM model, [EN22] showed that it can be implemented in the LOCAL model through an $O(\operatorname{log}n/\rho)$-round randomized LOCAL algorithm.
Now let two nodes $v,v^{\prime}$, such that $\operatorname{dist}(v,v^{\prime})\leq k$. Let $u_{v}$ be the center of the cluster that $v$ joins, that is the node realizing $\min_{u\in V}\operatorname{dist}_{\text{shift}}(u,v)$. Let $d_{1}\leq\ldots\leq d_{n}$ be the distances from $v_{1}$ to any node $v_{1}=u_{1},\ldots,u_{n}\in V$. Consider $\lambda_{i}$ to be the exponential random variable associated to the node $u_{i}$. By Lemma 3.5, we have that, with probability $O(\rho k/2)$, it holds that the smallest value and the second-smallest value in the family $\{(d_{i}-\lambda_{i})\}_{i\in[n]}$ are within distance $k$ of each other. Let $k=1/(c\rho)$ for a large enough constant $c$ so that the probability $O(\rho k/2)$ is strictly less than $1/2$.
[…]
3.2 Algorithm description
Our algorithm simply repeats the following process $\operatorname{log}n$ times: Compute a clustering, such that for every node $v$ with constant probability its entire $O(\operatorname{log}^{4}n)$ neighborhood is contained within its cluster. In each cluster we brute force a maximal sequence of improving sets, each of improving ratio at least $R$. We then apply the relabeling implied by this sequence in every cluster and inform all nodes of the cluster. Informally, we now consider the nodes in these clusters to be ”safe”. To ensure that we do not ”undo” the progress we have made in previous iterations, we have to increase the minimum improving ratio $R$ between every phase by a small value (roughly $\varepsilon$).
Algorithm 1 LOP algorithm (from the perspective of a node)
1:Number of nodes $n$, description of the LOP $(\Pi,\Psi)$.
2:Solution to $(\Pi,\Psi)$.
3:Initialize $\lambda\leftarrow$ minimum improvement associated to $(\Pi,\Psi)$
4:Initialize $R\leftarrow\lambda/4$
5:Initialize $\varepsilon\leftarrow\frac{\lambda}{100c_{1}\operatorname{log}n}$ $\triangleright$ $c_{1}$ is a constant given by the proof of Theorem 3.7
6:Initialize any output label for the node and its incident half-edges
7:The next loop identifies the “phases” of our algorithm. In the following, $c$, $c_{1}$, $c_{2}$, and $c_{3}$ are large enough constants. The constants $c_{2}$ and $c_{3}$ are given by Lemmas 4.5 and 4.8, while $c_{1}$ is given by the proof of Theorem 3.7, and $c$ is given by Lemma 3.6.
8:for $i=1$ to $c_{1}\operatorname{log}n$ do
9:  Run MPX with $\rho=10\cdot c_{2}\cdot\varepsilon^{2}/(c\operatorname{log}^{2}n)$ $\triangleright$ $d=O(\operatorname{log}n/\rho)$ by Lemma 3.6
10:  $C\leftarrow$ cluster of the node
11:  $H\leftarrow$ graph induced by $C$
12:  $\ell\leftarrow$ labeling of $H$
13:  if the node is the cluster leader then
14:   Choose any maximal sequence of $R$-improving sets $\left((A_{i},\ell_{i})\right)_{i=1}^{k}$ in $H$ w.r.t. $\ell$ such that
15:      1. $\mathcal{N}_{2r+1}[A_{i}]\subseteq C$ for all $i\in[k]$
16:      2. The diameter of each $A_{i}$ is at most $c_{3}\operatorname{log}(n)/\varepsilon$
17:   Adopt the new labeling according to $\ell_{k}$
18:   $\ell\leftarrow\ell_{k}$
19:   Broadcast $\ell_{k}$ to all nodes in $C$
20:  else
21:   Wait for the broadcast by the leader
22:   Adopt the new labeling according to $\ell_{k}$
23:   $\ell\leftarrow\ell_{k}$
24:  end if
25:  Set $R\leftarrow R+\frac{\lambda}{20c_{1}\operatorname{log}n}$
26:end for
27:return labels of the node and of the incident half-edges
We now give a formal description of the algorithm given in Algorithm 1. It solves an LOP $(\Pi,\Psi)$ as defined in Definition 2.7. It takes as input the number of nodes $n$ and the description of the LOP $(\Pi,\Psi)$. In more detail, it gets the full description of $\Pi=(\mathcal{V}_{\operatorname{in}},\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}},\mathcal{E}_{\operatorname{out}},\mathcal{C})$, where $\mathcal{C}$ is an $(2r,\Delta)$-set of constraints over $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$, and the potential function $\Psi$ that assigns a non-negative real value to every element of $\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},r,\Delta)$.
For every labeled centered graph $(H,v_{H})\in\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},2r,\Delta)$ that does not belong to $\mathcal{C}$, there is a way to reassign the labels to $v_{H}$ and its incident half-edges so that the potential $\operatorname{Pot}(G,\ell)$ decreases and the centered graph $(H,v_{H})$ with the new labeling belongs to $\mathcal{C}$. Let $\ell^{\prime}$ be the new labeling: $(\{v_{H}\},\ell^{\prime})$ is a minimal improving set with some positive improvement $\hat{\lambda}$. Let $\lambda$ be the minimum over all possible values of $\hat{\lambda}$. We say that $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$ (which is initialized in Line 3). Similarly, let $\Lambda$ be the maximum possible decrease of the potential over any centered graph $(H,v_{H})\in\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},2r,\Delta)$. We say that $\Lambda$ is the maximum improvement associated to $(\Pi,\Psi)$.
The idea of the algorithm is as follows. Set $\varepsilon=\lambda/(100c_{1}\operatorname{log}(n))$, where $c_{1}$ is a small enough constant (Line 5), and set a threshold improving ratio to be $R=\lambda/4$ (Line 4). We start with an arbitrary labeling of the nodes and half-edges of $G$. Then, we have $\Theta(\operatorname{log}n)$ phases (from Line 8 to Line 26), where in each phase we run the MPX clustering algorithm (Line 9) with parameters $\rho=10\cdot c_{2}\cdot\varepsilon^{2}/(c\operatorname{log}^{2}n)=O(\operatorname{log}^{4}n)$ and $d=O(\operatorname{log}n/\rho)$, where $c,c_{2}$ are large enough constants that come from Lemmas 3.6 and 4.8, respectively: the bound on $d$ then holds by Lemma 3.6. Each run of MPX takes $O(d)=O(\operatorname{log}^{5}n)$ rounds. We assume that each cluster $C$ given by the MPX algorithm comes with a leader node $v_{C}$, that can be the center of the cluster itself. This anyway can be obtained in $O(d)$ rounds after each MPX run. Each leader node looks at the subgraph $H$ of $G$ induced by the cluster $C$ it belongs to, and it computes a maximal sequence of $R$-improving sets $(A_{i},\ell_{i})_{i=1}^{k}$ in $H$ w.r.t. $\ell$. The sequence satisfies that $\mathcal{N}_{2r+1}[A_{i}]\subseteq C$ for all $i\in[k]$ (from Line 13 to Line 14), ensuring that the sequences of different clusters do not interfere with each other. Furthermore, we require the diameter of each set in the sequence to be at most $c_{3}\operatorname{log}n/\varepsilon$ (the choice of the large constant $c_{3}$ comes from Lemma 4.5). When we say that the sequence is maximal, we mean that in $H$ labeled by $\ell_{k}$ there is no minimal improving set $(A^{\prime},\ell^{\prime})$ such that $\operatorname{IR}(A^{\prime},\ell_{k},\ell^{\prime})>R$ and $\mathcal{N}_{2r+1}[A^{\prime}]\subseteq H$. Note that such a sequence must exist and terminate, as the potential of $H$ is always reduced after every relabeling (and cannot be reduced arbitrarily): furthermore, the sequence can be computed in $O(d)$ rounds by brute-force. For this the leader simply collects the topology and labeling of its entire cluster and brute forces every possible sequence in local computation. Then, $v_{C}$ adopts the labeling defined by $\ell_{k}$ and also broadcasts $\ell_{k}$ to every node of the cluster $C$ (Line 17). Every non-leader node just waits for the leader node to broadcast the new labeling (lines 20 and 21), and then it adopts the new labeling (Line 22). Then, the phase terminates and the nodes update the reference value of the improving ratio $R$ to $R+\lambda/(20c_{1}\operatorname{log}n)$ (Line 25). The algorithm terminates after $\Theta(\operatorname{log}n)$ phases, and it returns the labels of the nodes and of the incident half-edges (Line 27).
Note that the round complexity of the algorithm is the running time of each phase, that is, $O(d)$, multiplied by the number of phases, that is, $O(\operatorname{log}n)$. Hence,
$T(n)=O(\operatorname{log}n\cdot d)=O(\operatorname{log}^{4}n/\varepsilon^{2})=O(\operatorname{log}^{6}n).$
In the rest of the paper, we prove the following theorem.
Theorem 3.7.
Let $(\Pi,\Psi)$ be a LOP as defined in Definition 2.7. The randomized LOCAL algorithm described in Algorithm 1 solves $(\Pi,\Psi)$ in $O(\operatorname{log}^{6}n)$ rounds w.h.p.
4 Analysis of the algorithm
In this section, we prove Theorem 3.7. We already analyzed the runtime of the algorithm presented in Section 3.2. Hence, in the following, we prove that the algorithm produces a proper solution to the given problem.
For the rest of the section, we fix an LOP of interest $(\Pi,\Psi)$ and any input graph $G=(V,E)$ of $n$ nodes that is already $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled (hence it has an input and a – possibly invalid – output labeling). Let $\ell$ denote the labeling of $G$. Recall that $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$, while $\Lambda$ is the maximum improvement associated to $(\Pi,\Psi)$. Trivially, $\lambda\leq\Lambda$. We remark that all the values $r,\Delta,\lambda,\Lambda$ are constant values and will be hidden in the asymptotic notation. In the proofs of the lemmas in Section 4.1, we do keep track of a loose dependency in these parameters, which is polynomial. The reader might extract a $\text{poly}(\Delta^{r},\Lambda)\cdot\text{polylog}(n)$-time algorithm from this analysis; however, we remark that we did not optimize such dependencies.
We fix some standard notation.
Definition 4.1.
For any subset of nodes $A\subseteq V$, we denote by $\ell^{\star}_{A}$ the best possible relabeling of the nodes of $A$ and their incident half-edges, that is,
$\ell^{\star}_{A}=\arg\max_{x\in\mathscr{L}}\set{\operatorname{Imp}(A,\ell,x)}$,
breaking ties arbitrarily, where $\mathscr{L}$ is the set of all possible relabelings of $G$ such that $G$ is a $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph where the input labels are untouched, and the output labels might change only on the nodes of $A$ and their incident half-edges.
The whole analysis of the algorithm relies on the following two lemmas (Lemmas 4.2 and 4.4), the last of which requires involved technical claims. The first lemma states that each phases terminates.
Lemma 4.2.
For any phase $j$ of the algorithm, let $\ell_{j}$ be the labeling of the algorithm after finishing phase j1j-1 (Lines 1823) and $R_{j}$ be the value of $R$ at the end of phase j1j-1 (Line 25). Furthermore, let $C^{(j)}_{1},\ldots,C^{(j)}_{k}$ be the clusters that are the result of the MPX algorithm at phase $j$ (Line 9), Then, for each cluster $C^{(j)}_{i}$, any maximal sequence of $R_{j}$-improving sets is finite.
Proof.
Let $H^{(j)}_{i}$ be the graph induced by cluster $C^{(j)}_{i}$, and $\ell_{H^{(j)}_{i}}$ the restriction of $\ell_{j}$ to $H^{(j)}_{i}$. Since $\operatorname{Pot}(H^{(j)}_{i},\ell_{H^{(j)}_{i}})$ is a finite value, and each improving set decreases this value by at least some constant $R_{j}>0$, then the sequence cannot have more than $\operatorname{Pot}(H^{(j)}_{i},\ell_{H^{(j)}_{i}})/R_{j}$ elements. ∎
The second lemma states that, at the end of some phase $i$, all improving sets are “close enough” to the borders of the current and previously computed clusters. Let us be more formal.
Definition 4.3 (Border sets).
For phase $i$ of the algorithm, we define
$B_{i}=\left\{u\in V\ \middle|\ \exists\{u,v\}\in E,C_{u}\neq C_{v}\right\}$
where $C_{u}$ is the cluster of $u$ and $C_{v}$ the cluster of $v$ after running MPX in phase $i$ (Line 9).
So $B_{i}$ is the set of all nodes that have an edge that is crossing different clusters after running MPX at phase $i$. The next lemma shows that after phase $i$, all remaining improving sets of small diameter that have improving ratio at least $R_{i}$ (the value of $R$ initialized at the end of phase i1i-1 at Line 25) are close to all borders $B_{j}$ for all $j\leq i$. Note that after the last phase the intersection of all border nodes will be empty with high probability. So this lemma implies that at the end of our algorithm, there do not exist any improving sets with small diameter and good improving ratio anymore.
Lemma 4.4.
At the end of phase $i$ of the algorithm, all minimal improving sets of diameter at most $c_{3}\operatorname{log}n/\varepsilon$ and improving ratio at least $R_{i}$ (the value of $R$ initialized at the end of phase i1i-1 at Line 25) are fully contained in the set
$\mathcal{N}_{t_{1}}[B_{i}]\cap\left(\bigcap_{j=1}^{i-1}\mathcal{N}_{t_{2}}[B_{j}]\right),$
for $t_{1}=c_{3}\operatorname{log}n/\varepsilon+10(r+1)$ and $t_{2}=c_{2}\operatorname{log}^{2}n/\varepsilon^{2}+t_{1}$, where $c_{2}$ and $c_{3}$ are as in Lines 9 and 16, respectively.
The proof of Lemma 4.4 relies on the following two lemmas. The first one states that for any minimal improving set $A$ of some improving ratio $\beta$, for any node $v\in A$, we can find, within some radius-$O(\operatorname{log}n/\varepsilon)$ neighborhood of $v$, a minimal improving set $A^{\prime}\subseteq A$ with improving ratio at least $\beta-\varepsilon$. The proof of Lemma 4.5 is involved, and we defer it to Section 4.1.
Lemma 4.5.
Let $\ell$ be a labeling of $G=(V,E)$ and let $(S,\ell_{S})$ be a $\beta$-minimal improving set, for some $\beta>0$. Recall that $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$. Then, for every $0<\varepsilon<\min{\{\beta,\lambda\}}$ and any $v\in S$ there exists a large enough constant $c_{3}$, a $k\leq c_{3}\operatorname{log}n/\varepsilon$, and a subset $A\subseteq S\cap\mathcal{N}_{k}[v]$, such that $\operatorname{IR}(A,\ell,\ell_{A}^{\star})\geq\beta-\varepsilon$. Furthermore, $(A,\ell_{A}^{\star})$ is minimal.
The second lemma states that in a sequence of $\beta$-improving sets with small diameter, before any relabeling, around each node of the sequence there is a minimal improving set $A$ with improving ratio slightly less than $\beta$. Importantly, this $A$ is an improving set w.r.t. the original labeling, before doing any relabeling of the sequence. If, after the sequence, we have an improving set $A$ in some area which we previously already processed in phase $j$, then this lemma implies that we would have already seen a minimal improving set $A^{\prime}$ in this same area in phase $j$. Because $A^{\prime}$ will have a slightly worse improving ratio than $A$, we have to increase the improving ratio $R$ between every phase to make this argument useful. The proof of Lemma 4.6 is also quite involved, and we defer it to Section 4.1.
Lemma 4.6.
Let $(A_{i},\ell_{i})_{i=1}^{h}$ be a sequence of $\beta$-improving sets of diameter at most $c_{3}\operatorname{log}n/\varepsilon$, for some $\beta>0$ and $0<\varepsilon<\min{\{\beta,\lambda\}}$, where $c_{3}$ is the constant given by Lemma 4.5, and $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$. Then, for each $i\in[h]$ and each $v\in A_{i}$, there exists a large enough constant $c_{2}$ and an improving set $(A,\ell_{A})$ of diameter $c_{3}\operatorname{log}n/\varepsilon$, such that $\operatorname{IR}(A,\ell,\ell_{A})\geq\beta-2\varepsilon$, that is fully contained in the set $\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[v]\cap(\cup_{j\in[h]}A_{j})$. Furthermore, $(A,\ell_{A})$ is minimal.
We first proceed with the proof of Lemma 4.4 and then with that of Theorem 3.7.
Proof of Lemma 4.4.
For any phase $i>0$, let $\mathcal{C}^{(i)}_{1},\ldots,\mathcal{C}^{(i)}_{h_{i}}$ be the clusters determined by MPX at phase $i$ (Line 9), let $\ell_{i}$ be the labeling of $G$ at the end of phase $i$ (Lines-18 and 23), and let $R_{i}$ be the improving ratio initialized at the end of phase i1i-1 (Line 25), with $R_{1}=R$ as in Line 4. We proceed by induction on the phase $i$.
If $i=1$, $R_{1}=R$ as initialized at Line 4. A minimal improving set $(A,\ell_{A})$ of diameter at most $c_{3}\operatorname{log}n/\varepsilon$ with $\operatorname{IR}(A,\ell_{1},\ell_{A})\geq R_{1}$ must be such that there is no $s\in[h_{1}]$ such that $\mathcal{N}_{2r+1}[A]\subseteq C^{(1)}_{s}$, otherwise we are breaking maximality in Line 14. Hence, $\operatorname{dist}(A,B_{1})\leq 2r$ and we have that $A\subseteq\mathcal{N}_{c_{3}\operatorname{log}n/\varepsilon+2r}[B_{1}]$, as the diameter of $A$ is at most $c_{3}\operatorname{log}n/\varepsilon$. Now, assume $i>1$ and the statement to be true for all phases $j=1,\ldots,i-1$. Let $(A,\ell_{A})$ be a minimal improving set of diameter at most $c_{3}\operatorname{log}n/\varepsilon$ such that $\operatorname{IR}(A,\ell_{i},\ell_{A})\geq R_{i}$. For the same reason as in the case $i=1$, it holds that $A\subseteq\mathcal{N}_{c_{3}\operatorname{log}n/\varepsilon+2r}[B_{i}]$. Suppose, by contradiction, that $A$ is not contained in
$\bigcap_{j=1}^{i-1}\mathcal{N}_{t_{2}}[B_{j}].$
This implies that there is a $j^{\star}\leq i-1$ such that $A$ is not contained in $\mathcal{N}_{t_{2}}[B_{j^{\star}}]$. Note that the diameter of $A$ is at most $c_{3}\operatorname{log}n/\varepsilon$ and at least one node of $A$ is not in $\mathcal{N}_{t_{2}}[B_{j^{\star}}]$. As a result, the set $\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}+2r+1}[A]$ is fully contained in some cluster $C^{(j^{\star})}_{s}$, for some $s\in[h_{j^{\star}}]$.
We can now describe the family of all improving sets that are actually used by our algorithm after phase $j^{\star}$ (Line 14) as a single improving sequence. We will define an ordering of all minimal improving sets that our algorithm has flipped. In order to do that, we first define an ordering of all clusters. We say that $C^{(m_{1})}_{s_{1}}<C^{(m_{2})}_{s_{2}}$ if and only if:
  • $m_{1}<m_{2}$, or
  • $m_{1}=m_{2}$ and $s_{1}<s_{2}$.
In each cluster $C^{(m)}_{s}$ we also consider the natural ordering of the sets composing the sequence $(S^{(m)}_{s,t},\ell_{S^{(m)}_{s,t}})_{t\in[k_{s}^{(m)}]}$ of $R_{m}$-improving sets that were chosen by the leader of $C^{(m)}_{s}$ (Line 14), where $k_{s}^{(m)}$ is the number of minimal improving sets that have been flipped inside $C^{(m)}_{s}$ at phase $m$. We can create a sequence of $R_{j^{\star}+1}$-improving sets by concatenating all these maximal sequences
$\left(\left((S^{(m)}_{s,t},\ell_{S^{(m)}_{s,t}})_{t\in[k_{s}^{(m)}]}\right)_{s\in[h_{m}]}\right)_{j^{\star}+1\leq m\leq i}$
and at the end we put the improving set $(A,\ell_{A})$. Observe that $(A,\ell_{A})$ is such that $\operatorname{IR}(A,\ell_{i},\ell_{A})\geq R_{i}\geq R_{j^{\star}+1}$ and hence the whole sequence is an $R_{j^{\star}+1}$-improving sequence. Let us rename this sequence as $(\hat{S}_{t})_{t\in[t_{i}]}$, where $t_{i}$ is the overall length of the defined sequence.
Now, for each $v\in A$, by Lemma 4.6 we have that there exists a minimal improving set $(A^{\prime},\ell_{A^{\prime}})$ that is fully contained in the set
$\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[v]\cap\left(\cup_{t\in[t_{i}]}\hat{S}_{t}\right),$
and such that $\operatorname{IR}(A^{\prime},\ell_{j^{\star}},\ell_{A^{\prime}})\geq R_{j^{\star}+1}-2\varepsilon$. By Lemma 4.5, for each $v\in A^{\prime}$, there exists a minimal improving set $(A^{\prime\prime},\ell_{A^{\prime\prime}})$ such that $A^{\prime\prime}\subseteq A^{\prime}\cap\mathcal{N}_{c_{3}\operatorname{log}n/\varepsilon}[v]$ and $\operatorname{IR}(A^{\prime\prime},\ell_{j^{\star}},\ell_{A^{\prime\prime}})\geq R_{j^{\star}+1}-3\varepsilon\geq R_{j^{\star}}$. The fact that
$A^{\prime\prime}\subseteq\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[v]$
implies that $A^{\prime\prime}\subseteq\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[A]$. Since $\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}+2r+1}[A]$ is fully contained in some cluster $C^{(j^{\star})}_{s}$, so it is $\mathcal{N}_{2r+1}[A^{\prime\prime}]$. Hence, the sequence $(S^{(j^{\star})}_{s,t},\ell_{S^{(j^{\star})}_{s,t}})_{t\in[k_{s}^{(j^{\star})}]}$ was not maximal, reaching a contradiction. ∎
With Lemmas 4.4 and 4.2 in hand, we can finally prove Theorem 3.7.
Proof of Theorem 3.7.
By Lemma 4.2, we have that the algorithm terminates. As discussed in Section 3.2, the running time of the algorithm is $O(\operatorname{log}^{6}n)$. We just have to argue about correctness.
Let $i=c_{1}\operatorname{log}n$ be the last phase, and let $\ell_{i}$ be the labeling obtained at the end of phase $i$ (Lines 18 and 23). By contradiction, suppose that there is an error in the graph, that is, a centered graph $(H,v_{H})$ of radius $2r$ that does not belong to $\mathcal{C}$. This means that $v_{H}$ can change the output labels of itself and its incident edges to get an improvement of at least $\lambda$. Let $\ell_{v}$ be that relabeling, so that we can get the improving set $(\set{v_{H}},\ell_{v})$, with $\operatorname{IR}(\set{v_{H}},\ell_{i},\ell_{v})\geq\lambda$. Let us now take the minimal version $(\set{v_{H}},\ell_{v}^{\star})$ of $(\set{v_{H}},\ell_{v})$.
Since $R_{c_{1}\operatorname{log}n}\leq\lambda$ by definition, Lemma 4.4 implies that $\set{v_{H}}$ is fully contained in
$\mathcal{N}_{t_{1}}[B_{i}]\cap\left(\bigcap_{j=1}^{i-1}\mathcal{N}_{t_{2}}[B_{j}]\right).$
Notice that, by Lemma 3.6, at each phase $j\leq i$, for every node $v$ there is probability at least $1/2$ that $\mathcal{N}_{t_{2}+1}[v]$ is fully contained within some cluster of the $j$-th run of MPX. Note that $(1/2)^{c_{1}\operatorname{log}n}=1/n^{c_{1}}$. Hence, with probability $1-1/n^{c_{1}}$, $\{v_{H}\}$ is not contained in
$\mathcal{N}_{t_{1}}[B_{i}]\cap\left(\bigcap_{j=1}^{i-1}\mathcal{N}_{t_{2}}[B_{j}]\right),$
reaching a contradiction with Lemma 4.4.
Now, by the union bound, with probability $1-1/n^{c_{1}-1}$, there is no node that is a center of a centered graph of radius $2r$ that is invalid according to $\mathcal{C}$, yielding the thesis. ∎
4.1 Technical claims and proofs
Here we prove Lemma 4.5 and Lemma 4.6. The first one basically says that, for any minimal improving set $A$ of some improving ratio $\beta$, for any node $v\in A$, we can find, within some radius-$O(\operatorname{log}n/\varepsilon)$ neighborhood of $v$, a minimal improving set $A^{\prime}\subseteq A$ with improving ratio at least $\beta-\varepsilon$. In order to prove it, we start with a preliminary lemma.
Lemma 4.7.
Consider any $S\subseteq V$, and let $B\subseteq S$. If $A=S\setminus B$, then
[…]
Recall that $\Lambda$ is the maximum improvement associated to $(\Pi,\Psi)$, and the labeling $\ell^{\star}_{M}$ is the best possible relabeling of $M$.
Proof.
We have that
$\operatorname{Imp}(S,\ell,\ell^{\star}_{S})=\operatorname{Pot}(G,\ell)-\operatorname{Pot}(G,\ell^{\star}_{S})$
[…]
where the last equality follows from the fact that the local potential of a graph centered at node $v$ of radius $r$ can only change if any label in the centered graph is modified, and hence $\Psi_{\ell}(G_{r}(v),v)=\Psi_{\ell^{\star}_{S}}(G_{r}(v),v)$ for all nodes at distance at least $r+1$ from $S$.
Let $\ell_{A}$ be the labeling defined as follows: On the nodes in $A$ and their incident half-edges, it is equal to $\ell^{\star}_{S}$, and on the rest of the graph it is equal to $\ell$. We similarly define $\ell_{B}$ (we just replace $B$ with $A$ in the previous definition).
We want to analyze the improvements of $(B,\ell_{B})$ and $(A,\ell_{A})$ compared to that of $(S,\ell^{\star}_{S})$. Consider any node $v$ such that its distance from $A$ (resp. $B$) is at least $r+1$. The two labelings $\ell_{A}$ (resp. $\ell_{B}$) and $\ell^{\star}_{S}$ will yield the same evaluation of the potential on the centered graph $(G_{r}(v),v)$.
Let $C=\mathcal{N}_{r}[A]\cap\mathcal{N}_{r}[B]$. The following holds:
$\operatorname{Imp}(B,\ell,\ell_{B})$
[…]
Notice that we can get the exact same equality when replacing $B$ with $A$:
$\operatorname{Imp}(A,\ell,\ell_{A})$
[…]
Bringing those together we recover the improvement for all graphs centered in nodes of $\mathcal{N}_{r}[S]\setminus C$ of radius $r$ when applying $\ell^{\star}_{S}$.
$\operatorname{Imp}(A,\ell,\ell^{\star}_{A})+\operatorname{Imp}(B,\ell,\ell^{\star}_{B})$
[…]
$+$ $\sum_{v\in\mathcal{N}_{r}[B]\setminus C}\left[\Psi_{\ell}(G_{r}(v),v)-\Psi_{\ell^{\star}_{S}}(G_{r}(v),v)\right]+\sum_{v\in C\cap\mathcal{N}_{r}[B]}\left[\Psi_{\ell}(G_{r}(v),v)-\Psi_{\ell_{B}}(G_{r}(v),v)\right]$
[…]
- $\sum_{v\in C}\left[\Psi_{\ell_{A}}(G_{r}(v),v)+\Psi_{\ell_{B}}(G_{r}(v),v)-2\Psi_{\ell}(G_{r}(v),v)\right]$
[…]
- $\sum_{v\in C}\left[\Psi_{\ell_{A}}(G_{r}(v),v)+\Psi_{\ell_{B}}(G_{r}(v),v)-2\Psi_{\ell}(G_{r}(v),v)\right]$
[…]
The claim follows by the definition of $C$ and by rearranging terms. ∎
With Lemma 4.7 in hand, we can now prove Lemma 4.5.
Proof of Lemma 4.5.
For any subset $A\subseteq S$ we again define $B=S\setminus A$. Let $x_{S}=\operatorname{IR}(S,\ell,\ell_{S}^{\star})$. Starting with the inequality of Lemma 4.7 and using that $(S,\ell^{\star}_{S})$ is minimal, we have
[…]
For the sake of readability, we set $y=2\Lambda$. Fix $v\in S$ and consider the following “candidate sets”: $C_{i}=\mathcal{N}_{i}[v]\cap S$ for $i\geq 0$. Let $|n_{i}=\absolutevalue{\mathcal{N}_{r}[C_{i}]}$, and let $k$ be the smallest integer $i\geq 2r$ such that $n_{i}\leq n_{i-2r}(1+\varepsilon/(y(\Delta^{r}+1)))$.
We now prove that such $k$ must exist and must be at most $4ry(\Delta^{r}+1)\ln n/\varepsilon+2r+1$. It holds that $n_{k-1}>n_{k-1-2r}(1+\varepsilon/(y(\Delta^{r}+1)))>\ldots>(1+\varepsilon/(y(\Delta^{r}+1)))^{\left\lfloor{(k-1)}/{2r}\right\rfloor}$. By contradiction, assume $k-1>4ry(\Delta^{r}+1)\ln n/\varepsilon+2r$. By the known inequality $(1+x/y)^{y}\geq\exp[xy/(x+y)]$ for all $x>0,y>0$, we have that
$(1+\varepsilon/(y(\Delta^{r}+1)))^{\left\lfloor(k-1)/2r\right\rfloor}$
$=\$$\left(1+\frac{\varepsilon}{y(\Delta^{r}+1)}\cdot\frac{\left\lfloor(k-1)/2r\right\rfloor}{\left\lfloor(k-1)/2r\right\rfloor}\right)^{\left\lfloor(k-1)/2r\right\rfloor}$
[…]
where the last inequality follows from the fact that $\varepsilon\leq y(\Delta^{r}+1)$ (as $\varepsilon<\lambda$, and $\lambda\leq\Lambda$). Since $(k-1)/2r>2y(\Delta^{r}+1)\ln n/\varepsilon+1$, we get that
$\exp\left[\frac{\varepsilon}{2y(\Delta^{r}+1)}\cdot\left\lfloor(k-1)/2r\right\rfloor\right]\geq\exp[\ln n]=n,$
and, hence, $n_{k-1}>n$, which is a contradiction.
If $C_{k}=S$, we are done. Assume that $C_{k}$ is strictly contained in $S$ and let $A=C_{k}$ and $B=S\setminus A$. We need to upper bound the size of $\mathcal{N}_{r}[A]\cap\mathcal{N}_{r}[B]$. We first show that $\mathcal{N}_{r}[A]\cap\mathcal{N}_{r}[B]\subseteq\mathcal{N}_{r}[C_{k}]\setminus\mathcal{N}_{r}[C_{k-2r}]$, and then we proceed bounding from above the cardinality of $\mathcal{N}_{r}[C_{k}]\setminus\mathcal{N}_{r}[C_{k-2r}]$ (which is, by definition, $n_{k}-n_{k-2r}$). Let $u\in\mathcal{N}_{r}[A]\cap\mathcal{N}_{r}[B]$. Trivially, $u\in\mathcal{N}_{r}[C_{k}]=\mathcal{N}_{r}[A]$. Furthermore, since $u\in\mathcal{N}_{r}[B]=\mathcal{N}_{r}[S\setminus C_{k}]$, there exists a node $w\in S\setminus C_{k}$ such that $\operatorname{dist}(u,w)\leq r$. Since $w\notin C_{k}$ but $w\in S$, we have that $\operatorname{dist}(v,w)\geq k+1$. By the triangle inequality, we have that
$\operatorname{dist}(w,v)\leq\operatorname{dist}(u,w)+\operatorname{dist}(u,v),$
which implies that
$\operatorname{dist}(u,v)\geq\operatorname{dist}(w,v)-\operatorname{dist}(u,w)\geq k+1-r.$
[…]
Now we want to prove Lemma 4.6. The key ingredient for Lemma 4.6 is the next lemma, which states the following: in a sequence of $\beta$-improving sets in $G$ w.r.t. some labeling $\ell$, for any node $v$ in any of the improving sets, there exists an improving set $(A,\ell_{A})$ in the radius-$O(\operatorname{log}^{2}n/\varepsilon^{2})$ neighborhood of $v$ (that is fully contained in the union of the improving sets) that has improving ratio at least $\beta-\varepsilon$ w.r.t. $\ell$. Now, we can take the minimal improving set $(A^{\prime},\ell_{A^{\prime}}^{\star})$ with $A^{\prime}\subseteq A$ and, by Lemma 4.5, we can find a set $A^{\prime\prime}\subseteq A^{\prime}$ such that $(A^{\prime\prime},\ell^{\star}_{A^{\prime\prime}})$ is a minimal improving set with improving ratio at least $\beta-2\varepsilon$ w.r.t. $\ell$ and diameter at most $O(\operatorname{log}n/\varepsilon)$. Hence, whenever there is a sequence of $\beta$-improving sets, before any relabeling, any node in the sequence can locally find small diameter improving sets with improving ratio at least $\beta-2\varepsilon$ w.r.t. $\ell$.
Lemma 4.8.
Let $0<\varepsilon<\min\{\beta,\lambda\}$ be any value, for some $\beta>0$, where $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$. Let $((A_{i},\ell_{i}))_{1\leq i\leq h}$ be a sequence of $\beta$-improving sets w.r.t. $\ell$ in $G$, where each minimal improving set has diameter at most $c_{3}\operatorname{log}n/\varepsilon$ ($c_{3}$ as in Line 16). Then for all $1\leq j\leq h$, for all $v\in A_{j}$ there exists a large enough constant $c_{2}>0$ and some $k\leq c_{2}\operatorname{log}^{2}n/\varepsilon^{2}$, such that there exists a set $A\subseteq\mathcal{N}_{k}[v]\cap(\bigcup_{1\leq i\leq h}A_{i})$ with $\operatorname{IR}(A,\ell,\ell^{*}_{A})\geq\beta-\varepsilon$.
Before proving Lemma 4.8, we remark that $c_{2}$ (Line 9 of the algorithm) is given in the statement of Lemma 4.8.
Proof.
We will employ an argument that is similar to that of Lemma 4.7, but this time we will work on an auxiliary graph. We define the auxiliary graph $H=(V_{H},E_{H})$ where $V_{H}$ contains a node $\sigma(A_{i})$ for each $A_{i}$, $i\in[h]$, and $\{\sigma(A_{i}),\sigma(A_{j})\}\in E_{H}$ if and only if $\operatorname{dist}_{G}(A_{i},A_{j})\leq 2r$.
The reason for defining the edges like this is the following: if the distance of $A_{i}$ and $A_{j}$ is at most $2r$, there are centered graphs of radius $r$ of which both the relabeling of $A_{i}$ and $A_{j}$ might affect the potential. Conversely, if $\sigma(A_{i})$ and $\sigma(A_{j})$ are not adjacent in $H$, then no node in $G$ can be affected by relabelings of both $A_{i}$ and $A_{j}$.
We define a weight function $w\colon V_{H}\to\mathbb{N}$ for the nodes in $H$ by $|w(x)=\absolutevalue{\sigma^{-1}(x)}$, that is, to any element $x=\sigma(A_{i})$ (for some $i\in[h]$), we assign a weight that is equal to the cardinality of $A_{i}$. With an abuse of notation, we define $w(B)=\sum_{x\in B}w(x)$ for every subset $B$ of $V_{H}$.
Observe that the improvement of the entire sequence of $\beta$-improving sets $((A_{i},\ell_{i}))_{1\leq i\leq h}$ is at least $\beta\cdot w(V_{H})$, due to the definition of sequence of $\beta$-improving sets. Now since the maximum improvement in the potential of any labeling in $G$ is at most $\Lambda\cdot n$, the overall improvement of the sequence itself cannot be larger than $\Lambda\cdot n$. This implies that $w(V_{H})\leq\Lambda\cdot n/\beta$.
Fix any $j\in[h]$, and consider an arbitrary $v\in A_{j}$. Let $x=\sigma(A_{j})\in V_{H}$. We define the “candidate sets” with regard to $z$ as $C_{i}=\mathcal{N}_{i}[z]$ for $i\geq 0$. Let $k$ be the minimum integer $i$ such that $\frac{\varepsilon}{K}w(C_{i})>w(C_{i+1}\setminus C_{i})=w(C_{i+1})-w(C_{i})$, for $K=(\beta-\varepsilon)+(\Delta^{r}+1)\Lambda$. We claim that $k\leq\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$. By contradiction, suppose that $k>\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$. Let $j=\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$. Then, we have that $w(C_{i+1})-w(C_{i})\geq\frac{\varepsilon}{K}w(C_{i})$ for all $i\leq j$. Then,
$w(C_{j})\geq(1+\varepsilon/K)w(C_{j-1})\geq(1+\varepsilon/K)^{3K\operatorname{log}(\Lambda n/\beta)/\varepsilon}.$
Again, by exploiting that $(1+x/y)^{y}\geq\exp[xy/(x+y)]$ for all positive $x,y$, and that $\varepsilon<\lambda\leq\Lambda$, we get that
$w(C_{j})\geq\exp\left[\frac{KM\operatorname{log}(\Lambda n/\beta)}{\varepsilon+K}\right]\geq(\Lambda n/\beta)^{3/2}.$
This is a contradiction since $w(C_{j})\leq w(V_{H})\leq\Lambda n/\beta$.
Let
$A$ $=\{i\in[h]\mid A_{i}=\sigma^{-1}(x)\text{ for some }x\in C_{k+1}\},\text{ and}$
$A^{\prime}$ $=\{i\in[h]\mid A_{i}=\sigma^{-1}(x)\text{ for some }x\in C_{k}\}.$
Furthermore, let $s=\max(A)$. We also define $\hat{A}=\cup_{i\in A}A_{i}$, and $\hat{A}^{\prime}=\cup_{i\in A^{\prime}}A_{i}$. Denote by $\ell_{A}$ the labeling obtained by setting $\ell_{A}=\ell_{s}$ for all nodes in $\hat{A}$ and their incident half-edges, and $\ell_{A}=\ell$ for all other nodes and half-edges.
Let $i_{1},\ldots,i_{t}$ denote the elements of $A$, where $i_{1}<\ldots<i_{t}=s$. We define $f_{i_{0}}=\ell$ and then a new relabeling $f_{i_{j}}$ as the labeling that outputs $\ell_{i_{j}}$ on all the nodes in $A_{i_{j}}$ and their incident half-edges, but outputs $f_{i_{j-1}}$ everywhere else. Now we can obtain $\ell_{A}$ by sequentially applying the relabelings $f_{i_{1}},f_{i_{2}},\ldots,f_{i_{t}}=\ell_{A}$.
Importantly, for each $j\in[t]$, by the definition of $f_{i_{j}}$ and $E_{H}$, the labels of all nodes in $\mathcal{N}_{2r}[\hat{A}^{\prime}]$ and their incident half-edges given by $f_{i_{j}}$ are exactly those given by $\ell_{i_{j}}$: in fact, if the labeling of a node $u\in\mathcal{N}_{2r}[\hat{A}^{\prime}]$ (or its incident half-edges) is modified by $\ell_{i_{j}}$, then $u\in A_{i_{j}}$ and $\operatorname{dist}_{G}(A_{i_{j}},\hat{A}^{\prime})\leq 2r$, implying that $i_{j}\in A$.
As a result, if $i_{j}\in A^{\prime}$, then
$|.\operatorname{Imp}(A_{i_{j}},f_{i_{j-1}},f_{i_{j}})=\operatorname{Imp}(A_{i_{j}},\ell_{i_{j-1}},\ell_{i_{j}})\geq\operatorname{Imp}(A_{i_{j}},\ell_{i_{j}-1},\ell_{i_{j}})\geq\beta\cdot\absolutevalue{A_{i_{j}}}$.
[…]
$\operatorname{Imp}(\hat{A},\ell,\ell_{A})=\$$\sum_{j=1}^{t}\operatorname{Imp}(A_{i_{j}},f_{i_{j-1}},f_{i_{j}})$
[…]
We now show that, by the definition of $K$, it holds that
$w(C_{k})\left(\beta-\frac{\varepsilon(\Delta^{r}+1)\Lambda}{K}\right)\geq w(C_{k+1})\left(\beta-\varepsilon\right).$
Recall that $w(C_{k+1})\leq w(C_{k})(1+\varepsilon/K)$. Hence,
$w(C_{k})\left(\beta-\frac{\varepsilon(\Delta^{r}+1)\Lambda}{K}\right)$ $\geq w(C_{k+1})\cdot\frac{K}{K+\varepsilon}\cdot\left(\beta-\frac{\varepsilon(\Delta^{r}+1)\Lambda}{K}\right)$
$=w(C_{k+1})\cdot\frac{\beta K-\varepsilon(K-(\beta-\varepsilon))}{K+\varepsilon}$
$=w(C_{k+1})\cdot\frac{K(\beta-\varepsilon)+\varepsilon(\beta-\varepsilon)}{K+\varepsilon}$
$=w(C_{k+1})\cdot\left(\beta-\varepsilon\right)\cdot\frac{K+\varepsilon}{K+\varepsilon}$
$=w(C_{k+1})\cdot\left(\beta-\varepsilon\right).$
Note that $C_{k+1}=\mathcal{N}_{k+1}[z]$. Thus, the weak diameter of $\hat{A}$ in $G$ is at most $(2k+1)\cdot(2r+c_{3}\operatorname{log}n/\varepsilon)$. In particular, since $k\leq\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$ and $K\leq\beta+(\Delta^{r}+1)\Lambda$, we have that the weak diameter of $\hat{A}$ in $G$ is at most $100((\beta+\Delta^{r}\Lambda)\cdot\operatorname{log}(\Lambda n/\beta)/\varepsilon+1)(r+c_{3}\operatorname{log}n/\varepsilon)$. Since by assumption $\Delta$ and $r$ are constants, we have that $\hat{A}\subseteq\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[v]$ for some large enough constant $c_{2}>0$. By noticing that $\ell^{\star}_{\hat{A}}$ can only increase the improving ratio w.r.t. $\ell_{{A}}$, we conclude the proof of Lemma 4.8. ∎
We conclude with the proof of Lemma 4.6.
Proof of Lemma 4.6.
We first apply Lemma 4.8, and we obtain an improving set $(A,\ell_{A})$ with improving ratio at least $\beta-\varepsilon$ that is fully contained in $\mathcal{N}_{c_{2}\operatorname{log}^{2}/\varepsilon^{2}}[v]\cap(\cup_{i\in[h]}A_{i})$. Then, we take the minimal version of it, that is, $(A^{\prime},\ell^{\star}_{A^{\prime}})$, with $A^{\prime}\subseteq A$, and we apply Lemma 4.5, which yields the thesis. ∎
Acknowledgments
This work was partially supported by the MUR (Italy) Department of Excellence 2023 - 2027 for GSSI, the European Union - NextGenerationEU under the Italian MUR National Innovation Ecosystem grant ECS00000041 - VITALITY – CUP: D13C21000430001, and the PNRR MIUR research project GAMING “Graph Algorithms and MinINg for Green agents” (PE0000013, CUP D13C24000430001). This project was also discussed at the Research Workshop on Distributed Algorithms (RW-DIST 2025) in Freiburg, Germany; we would like to thank all workshop participants and organizers for inspiring discussion.